OpenStudy (babynini):

Convert polar to rectangular coordinate r= 2/(1-cos theta)

OpenStudy (babynini):

\[r^2=(\frac{ 2 }{ 1-\cos \theta })r\]

OpenStudy (babynini):

\[r^2=(\frac{ 4r }{1r-rcos \theta) }\]

OpenStudy (babynini):

@freckles :)

OpenStudy (babynini):

@jim_thompson5910 sorry..me again :P

zepdrix (zepdrix):

Eyyy

zepdrix (zepdrix):

I'm not quite sure what you did with the r there. hmmm

zepdrix (zepdrix):

We starting with this?\[\Large\rm r=\frac{2}{1-\cos \theta}\]

OpenStudy (babynini):

Yeah, and then I multiplied both sides by r to get r^2 but I think I distributed it wrong on the right side xP

zepdrix (zepdrix):

Yah darn ;c I guess I would recommend a different first step. Multiply both sides by (1-cos theta)

OpenStudy (babynini):

(oops idk where that four came from either..hah) ok r(1-cos theta) = 2

zepdrix (zepdrix):

distribute the r :)

OpenStudy (babynini):

r-rcos(theta)=2

zepdrix (zepdrix):

Mmm k, what can we do with the rcos(theta)?

OpenStudy (babynini):

r-x=2

zepdrix (zepdrix):

Ooo! ok that works out nicely. Let's add the x to the other side. r=x+2

zepdrix (zepdrix):

how bout that lonely r? Any ideas? :)

OpenStudy (babynini):

square it! ..? :P

zepdrix (zepdrix):

So you want to square it because you're probably used to using this identity,\[\Large\rm r^2=x^2+y^2\]But we actually do have a way of dealing with an r, if we square root this equation we get,\[\Large\rm r=\sqrt{x^2+y^2}\]Ya? And then we can avoid introducing another silly r.

OpenStudy (babynini):

ou

zepdrix (zepdrix):

\[\Large\rm \sqrt{x^2+y^2}=x+2\]That step make sense? :o

OpenStudy (babynini):

Yep yep

zepdrix (zepdrix):

Hmm what next? :) Any ideas?

OpenStudy (babynini):

umm can we simplify that into just x+y = x+2? or is that illegal?

zepdrix (zepdrix):

that is very illegal! :)

zepdrix (zepdrix):

To deal with the square root, we'll have to `square` both sides.

OpenStudy (babynini):

haha k I thought that sounded bad xD

OpenStudy (babynini):

x^2+y^2=x^2+4x+4

zepdrix (zepdrix):

Cooool, looks good.

OpenStudy (babynini):

...is that the end?

zepdrix (zepdrix):

No. Cancel some stuff out if you're able. If this is not a circle, then we'd like to get it in the form y=(x stuff).

OpenStudy (babynini):

y^2=4x+4?

zepdrix (zepdrix):

Ya that's a good step :) Hmm your y is squared...

zepdrix (zepdrix):

how can we deal with the squareeeee? :D

OpenStudy (babynini):

sq root?

zepdrix (zepdrix):

yeaaaaaa

zepdrix (zepdrix):

square rooting a square, something special happens... member?

OpenStudy (babynini):

oh gosh..err

OpenStudy (babynini):

wouldn't it just become y?

zepdrix (zepdrix):

Mmm close,\[\Large\rm \sqrt{a^2}=\pm a\]

zepdrix (zepdrix):

Here is what it looks like without the plus/minus. https://www.desmos.com/calculator/q6rbowsuio The red is the entire bowl opening to the right. You can see that the `positive root` only accounts for half of that bowl, ya? :)

zepdrix (zepdrix):

positive root is the blue

OpenStudy (babynini):

oo yes I see.

OpenStudy (babynini):

\[\pm (y)=\sqrt{4x+4}\]

OpenStudy (babynini):

so we have to figure out what quadrant it is in

zepdrix (zepdrix):

\[\Large\rm y^2=4x+4\]\[\Large\rm \pm y=\sqrt{4x+4}\]\[\Large\rm y=\pm\sqrt{4x+4}\]

zepdrix (zepdrix):

Nahhh we don't need to do any fancy business like that. Do we? Hmm. I mean, we need both the plus and the minus. But do you have some method for checking usually?

OpenStudy (babynini):

haha ok. Er usually my prof just decides on his own lol

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

Yayyy you did it! \c:/ Good Job Ms Miriam!

OpenStudy (babynini):

thank you! good job to you too! ey there's no place for putting plus or minus in the answer box :o

zepdrix (zepdrix):

is there place for a comma? because you could input: \(\Large\rm \sqrt{4x+4},\qquad-\sqrt{4x+4}\)

OpenStudy (babynini):

oh good idea. Well sometimes the answer is not so far into the equation. For example: I put the answer r= sq(7sec2(theta)) and they wanted (r^2(cos2theta))=7 ..a much less simplified version o.0

zepdrix (zepdrix):

yikes :O

zepdrix (zepdrix):

What is your profile picture? Are those ... light bulbs ... being used as planters or something? 0_o

OpenStudy (babynini):

Yeah D: it's sometimes a guessing game -.- so maybe they want y^2=4x+4 actually ?!

OpenStudy (babynini):

lolxD yes!!

zepdrix (zepdrix):

oo interesting :3

zepdrix (zepdrix):

Yah maybe that's what they want.

OpenStudy (babynini):

haha I think it's a good idea x)

OpenStudy (babynini):

Yep that's what it wanted :) thank you!

zepdrix (zepdrix):

oh cool c: